OpenStudy (anonymous):

Write an equation in point-slope form of the line with slope –8 that contains the point J(–3, 2).

OpenStudy (rulnick):

y-2 = -8 (x-3)

OpenStudy (anonymous):

(y - 2) = -8(x + 3)